Stock A Index Model regression Model 2% + 1.2 (Rm-Rf) R-Square 0.576 Residual Standard Deviation 11.00%...

Stock A
Index Model regression Model 2% + 1.2 (Rm-Rf)
R-Square 0.576
Residual Standard Deviation 11.00%
Standard Deviation of excess returns 22.00%
Rf 3%
Rm 8%
Calculate Trenor Measure for A:
Answers:

4.36%

5.22%

5.94%

6.67%
Answer #1

Ra -Rf = 2% + 1.2*(Rm -Rf) = 2% + 1.2*(8%-3%) = 8.00%

Treynor measure = (Ra-Rf)/beta = 8%/1.2 = 6.67% (Option D)